OpenStudy (anonymous):

find the general term sequence fora^n for the geometric system. a1=-5 and a2=10

OpenStudy (anonymous):

i tried to solve this equation a2=-5+(2-1)(10)

OpenStudy (cwrw238):

a geometric series is of the form a, ar , ar^2.. general term is ar^(n-1)

OpenStudy (anonymous):

well it has a and the n is at the bottom instead at the top a^2 is wrong i cannot find the correct button for a/n i think its a[n]

OpenStudy (anonymous):

for the term sequence a[n]

OpenStudy (cwrw238):

if a = a1 = -5 and a2 = -10 then can you figure out what r is?

OpenStudy (cwrw238):

* a2 = 10

OpenStudy (anonymous):

well bellow i dont know if i set the problem up to start off solving below

OpenStudy (anonymous):

i have a2=-5+(2-1)(10)

OpenStudy (cwrw238):

that is an arithmetic sequence general term a1 + (n - 1)d

OpenStudy (anonymous):

that is correct

OpenStudy (cwrw238):

its confusing because you said its geometric

OpenStudy (anonymous):

well the given set up is a[n]=a1+(n-1) d and it ask for general term of sequence

OpenStudy (cwrw238):

that is the general term for an AS a2 = -5 + (2 -1)d = -5 + d from this you can find the value of d , the common difference

OpenStudy (anonymous):

i think d is 10 i am not sure

OpenStudy (anonymous):

i got the answer a[n]=5

OpenStudy (cwrw238):

10 = -5 + d d = 15

OpenStudy (cwrw238):

a(n) = -5 + (n - 1)15

OpenStudy (anonymous):

ok so you added 10+5=15

OpenStudy (cwrw238):

so for example a5 = -5 + (5-1)15 = -5 + 60 = 55

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

a(n)=-5+(5-1)15

OpenStudy (anonymous):

oh i see you have answered it

OpenStudy (anonymous):

thank you for taking the time with me

OpenStudy (cwrw238):

a(n) = -5 + (n-1)15 is general form you just plug a value in for n - e.g. for 10th term you plug in n = 10

OpenStudy (cwrw238):

welcome